OpenStudy (anonymous):

What is the value of g[f(3)] for the functions f(x) = 2x + 3 and g(x) = x + 1?

OpenStudy (gorv):

@InsanelyChaotic f(3)=9 g(9)=9+1 =10

sam (.sam.):

g[f(3)] g(2x+3) (2x+3)+1 2x+4 understand?

sam (.sam.):

Oh and 2(3)+4 6+4=?

OpenStudy (anonymous):

10:)